OpenStudy (anonymous):

The sum of three consecutive odd integers is 18 less than five times the middle number. Find the three integers

OpenStudy (anonymous):

x + (x + 2) + (x + 4) = 5(x + 2) -18 3x + 6 = 5x +10 - 18 3x + 6 = 5x - 8 2x = 14 x = 7 7, 9, 11